# Vanishing criteria for Ceresa cycles

### v1

@@ 4.6. Identifying the complex torus $\mathrm{J}(\mathbb{V}^{\mu_{3}})$
Let $\mathcal{E}_{i}\rightarrow S$ be the relative elliptic curve with equation $y^{2}=x^{3}+{\rm disc}(f)^{i}$. The previous paragraph shows that $\mathrm{J}(\mathbb{W})$ is isomorphic to $\mathcal{E}_{i}({\rm\mathbb{C}})$ for some $i\in\{0,1,2,3,4,5\}$. We show that $i=1$, using our previous results on bielliptic Picard curves in [27]. Let $T\subset S$ be the closed subscheme where $b=0$, parametrizing even quartic polynomials $f=x^{4}+ax^{2}+c$. For $f\in T({\rm\mathbb{C}})$, the $\mu_{3}$-action on $\mathcal{C}_{f}$ extends to a $\mu_{6}$-action. A calculation shows ${\rm disc}|_{T}=16c(-a^{2}+4c)^{2}$. Applying the singular cohomology realization functor to [27, Theorem 5.1], the relative complex torus $\mathrm{J}((\mathbb{V}|_{T({\rm\mathbb{C}})})^{\mu_{6}})\rightarrow T({\rm\mathbb{C}})$ is isogenous to $\mathcal{E}_{1}({\rm\mathbb{C}})|_{T({\rm\mathbb{C}})}\rightarrow T({\rm\mathbb{C}})$. Since $\mathrm{J}((\mathbb{V}|_{T({\rm\mathbb{C}})})^{\mu_{6}})$ is a subtorus of $\mathrm{J}((\mathbb{V}|_{T({\rm\mathbb{C}})})^{\mu_{3}})$ of the same dimension, they must be equal, hence $\mathrm{J}(\mathbb{W})|_{T({\rm\mathbb{C}})}$ is isogenous to $\mathcal{E}_{1}({\rm\mathbb{C}})|_{T({\rm\mathbb{C}})}$ over $T({\rm\mathbb{C}})$. On the other hand, let $i\in\{0,1,2,3,4,5\}$ be such that $\mathrm{J}(\mathbb{W})\simeq\mathcal{E}_{i}({\rm\mathbb{C}})$. Then $\mathrm{J}(\mathbb{W})|_{T({\rm\mathbb{C}})}\simeq\mathcal{E}_{i}({\rm\mathbb{C}})|_{T({\rm\mathbb{C}})}$, hence $\mathcal{E}_{1}({\rm\mathbb{C}})|_{T({\rm\mathbb{C}})}$ is isogenous to $\mathcal{E}_{i}({\rm\mathbb{C}})|_{T({\rm\mathbb{C}})}$.
@@
Given $s=(a,b,c)\in S({\rm\mathbb{C}})$, consider the closed subscheme $T=\{(at,bt,ct^{2})\colon t\in{\rm\mathbb{A}}^{1}_{{\rm\mathbb{C}}}\}\cap S$ of $S$ and the restriction $\pi\colon\mathcal{E}|_{T}\rightarrow T$. The variety $\mathcal{E}|_{T}$ is an open subscheme of an elliptic surface with Weierstrass equation $y^{2}=x^{3}+t^{4}g(t)$, where $g(t)\in{\rm\mathbb{C}}[t]$ has degree $\leq 2$, using the formula (4.1). There exists a dense open $U\subset S$ such that for all $s\in U({\rm\mathbb{C}})$, $g(t)$ has two distinct nonzero roots. Let $s\in U({\rm\mathbb{C}})$. Using Tate’s algorithm, we see that the elliptic surface has a singular fiber above $t=0$ with Kodaira type $\text{\rm IV}^{*}$, singular fibers above the roots of $g(t)$ with Kodaira type II, and is smooth above $t=\infty$. The presence of a fiber of type II implies that $\mathcal{E}|_{T}\rightarrow T$ has no torsion sections [43, Lemma 7.8]. Moreover the Shioda–Tate formula [43, Theorem 6.3, Proposition 6.6 and §8.8] shows that the group of sections of $\mathcal{E}|_{T}\rightarrow T$ is free of rank $2$. Since $\mathcal{E}|_{T}$ receives an $\mathcal{O}$-action, its group of sections is free of rank $1$ over $\mathcal{O}$. We will now show that these facts can be used to prove the claims of the proposition by varying $s$ in $U({\rm\mathbb{C}})$.
@@ 4.7. Proof of the vanishing criterion in the Chow group
We claim that $\sigma$ is not a torsion section. If it were torsion, then ${\rm AJ}(\kappa_{\mathcal{C}})$ would be a torsion section of $\mathrm{J}(\mathbb{V}^{\mu_{3}})$, hence, by Proposition 4.12, $\kappa_{f}$ would be torsion for every $f\in S({\rm\mathbb{C}})$. This is not the case, since $\kappa_{f}$ is of infinite order if $f=x^{4}+x^{2}+1$ by [27, Corollary 2.9]. We conclude that $\sigma$ is not a torsion section.
@@
4.8. Families of Picard curves with torsion Ceresa cycle of arbitrarily large order
Theorem 4.2 shows that there are infinitely many plane quartic curves over ${\rm\mathbb{Q}}$ with torsion Ceresa cycle, since we may take $a=c=-12$. In fact, we can find explicit families of torsion Ceresa cycles of arbitrarily large order by analyzing the function $f\mapsto(I(f),J(f))$.
Let $C_{f}$ be a Picard curve defined by the polynomial $f(x)=x^{4}+ax^{2}+bx+c$ over a field $k$ with homogenization $F(x,z)=x^{4}+ax^{2}z^{2}+bxz^{3}+cz^{4}$. Associated to $C_{f}$ is the genus $1$ curve $D_{f}$ with equation $y^{2}=F(x,z)$. Viewing $D_{f}$ as an elliptic curve with origin $(1:1:0)$, it is isomorphic to the elliptic curve $y^{2}=x^{3}-I(f)x/3-J(f)/27$. The point $(1:-1:0)$ is sent to $(-2a/3,b)$ in this new model. Conversely, given a nonzero $k$-point $(\alpha,\beta)$ on the elliptic curve $E_{I,J}\colon y^{2}=x^{3}-Ix/3-J/27$, the quartic
(4.3) $f(x)=x^{4}-3\alpha x^{2}/2+\beta x+(I/12-3\alpha^{2}/16)$
has the property that $(I(f),J(f))=(I,J)$. It follows that if $(I,J)\in k$ is such that $4J^{3}-J^{2}\neq 0$, then $f\mapsto(-2a/3,b)$ induces a bijection between the set of polynomials $f=x^{4}+ax^{2}+bx+c\in k[x]$ satisfying $(I(f),J(f))=(I,J)$, and the set of nonzero elements of $E_{I,J}(k)$. Using this description, we can cook up explicit 1-parameter families of torsion Ceresa cycles of arbitrarily large order, as in the following corollary of Theorem 4.2:
Corollary 4.16.
For every integer $N$, there exists a number field $k=k(N)$ such that there exists infinitely many $\bar{\rm\mathbb{Q}}$-isomorphism classes Picard curves $C_{f}$ defined over $k$ with Ceresa cycle $\kappa_{f}$ of order at least $N$ in ${\rm CH}^{2}(J_{f,\bar{k}})$.
@@ 4.8. Families of Picard curves with torsion Ceresa cycle of arbitrarily large order
Let $M$ be an integer satisfying the conclusion of Theorem 4.2. Fix a nonzero $D\in{\rm\mathbb{Q}}$ and let $(I,J)\in\bar{{\rm\mathbb{Q}}}^{2}$ be a torsion point of order $M\cdot N$ on the elliptic curve $y^{2}=4x^{3}+D$. Let $k$ be the field of definition of $(I,J)$. For $t\in k$, define $\lambda(t)=t^{3}-It/3-J/27$, $(\alpha_{t},\beta_{t})=(t\lambda(t),\lambda(t)^{2})$ and the quartic $f_{t}(x)=x^{4}-3\alpha_{t}x^{2}/2+\beta_{t}x+(I/12-3\alpha_{t}^{2}/16)$ using Formula (4.3). Since $(\alpha_{t},\beta_{t})$ defines a point on the elliptic curve $E_{\lambda(t)^{2}I,\lambda(t)^{3}J}$, the quartic $f$ has the property that $(I({f_{t}}),J({f_{t}}))=(\lambda(t)^{2}I,\lambda(t)^{3}J)$. If $t\in k$ is such that $\lambda(t)\neq 0$, then $(\lambda(t)^{2}I,\lambda(t)^{3}J)$ is an $N$-torsion point on the elliptic curve $y^{2}=4x^{3}+\lambda_{t}^{6}D$. By Theorem 4.2, $\kappa_{f_{t}}$ is torsion in ${\rm CH}^{2}(J_{f_{t},\bar{k}})$ and its torsion order $\operatorname*{ord}(\kappa_{f_{t}})$ is divisible by $N$. When varying over $t\in k$ such that $\lambda(t)\neq 0$, the Picard curves $C_{f_{t}}$ cover infinitely many $\bar{\rm\mathbb{Q}}$-isomorphism classes, using [13, Lemma 1.21(b)] and the fact that $\beta_{t}^{2}/\alpha_{t}^{3}=\lambda(t)/t^{3}$ is not a constant function of $t$. ∎
Conversely, for any number field $k$, the order of a torsion Ceresa cycle $\kappa(C_{f})$, with $C_{f}$ defined over $k$, is bounded, with the bound depending only on the degree of $k$ over ${\rm\mathbb{Q}}$:
Corollary 4.17.
For every $d\geq 1$, there exists $N=N(d)\geq 1$ such that for every Picard curve $C_{f}$ over a number field $k$ of degree $d$, the order of $\kappa_{f}$ in ${\rm CH}_{1}(J_{f,\bar{k}})$ is either infinite or less than $N$.
@@
This follows from Theorem 4.2 and the uniform bound (depending only on $d$) on the order of a $k$-rational torsion point on an elliptic curve $y^{2}=x^{3}+D$ over any number field $k$ of degree $d$. ∎
5. Automorphism strata and Ceresa vanishing loci in genus $3$
@@
These vanishing loci have the following basic properties: $V_{g}^{\mathrm{rat}}\subset V_{g}^{\mathrm{alg}}\subset\mathcal{M}_{g}$; the hyperelliptic locus is contained in $V_{g}^{\mathrm{rat}}$; and $V_{g}^{\mathrm{alg}}\neq\mathcal{M}_{g}$ by Ceresa’s famous result [14]. It would be interesting to obtain further information about the components of $V_{g}^{\mathrm{rat}}$ and $V_{g}^{\mathrm{alg}}$. We end our paper by determining the automorphism group strata in $\mathcal{M}_{3}$ that are contained in $V_{3}^{\mathrm{rat}}$ or $V_{3}^{\mathrm{alg}}$.
So let $g=3$ and consider the open subscheme $\mathcal{M}_{3}^{\mathrm{nh}}\subset\mathcal{M}_{3}$ of non-hyperelliptic curves. There is a stratification $\mathcal{M}_{3}^{\mathrm{nh}}=\sqcup X_{G}$ into locally closed subvarieties such that a non-hyperelliptic curve $C$ over ${\rm\mathbb{C}}$ belongs to $X_{G}$ if and only if ${\rm Aut}(C)\simeq G$. It turns out that $X_{G}$ is irreducible and the closure of $X_{G}$ is a union of other strata. We refer to [31, §2.2] and references therein for a complete description of the loci $X_{G}$ and the closure relations between them. We reproduce here a diagram capturing these closure relations:
@@
For $n\in\{16,48,96\}$, the symbol $G_{n}$ the group of order $n$ and GAP label $(16,13),(48,33),$ and $(96,64)$ respectively [21]. See [31, Table 2] for models for a generic plane quartic in $X_{G}$. We make explicit the strata that are relevant for us: $\overline{X_{C_{3}}}$ is the locus of Picard curves studied in §4; $\overline{X_{C_{6}}}$ is the locus of bielliptic Picard curves studied in [28]; and the zero-dimensional strata each consist of a single automorphism-maximal curve with equation
@@
We end with the proof of Theorem D, which we restate for convenience:
@@
    Our analysis of Picard curves (Theorem C) shows that $X_{C_{3}}\subset V_{3}^{\mathrm{rat}}$, so $X_{G}\subset V_{3}^{\mathrm{rat}}$ for $G=C_{6},C_{9}$ and $G_{48}$ as well. On the other hand, the Ceresa cycle of the Fermat quartic and Klein quartic are known to be of infinite order in the Griffiths group; see [11, Theorem (4.1)] for the former and [24, §4] for the latter. By our observation, this means that $X_{G}\not\subset V_{3}^{\mathrm{rat}}$ for every stratum whose closure contains one of these curves. Since every stratum not contained in $\overline{X_{C_{3}}}$ has this property, we conclude the proof.
@@
    Since $V_{3}^{\mathrm{rat}}\subset V_{3}^{\mathrm{alg}}$, Part $(1)$ implies that $X_{G}\subset V_{3}^{\mathrm{rat}}$ only if $G=C_{3},C_{6},C_{9}$ or $G_{48}$. The criterion of Theorem A applies to the curves in $X_{C_{9}}$ and $X_{G_{48}}$ (see Example 3.1), so $V_{3}^{\mathrm{rat}}$ contains these strata. On the other hand, there exist curves in $X_{C_{6}}$ with nonvanishing $\kappa(C)$, by [27, Corollary 1.2]. So $X_{C_{6}}\not\subset V_{3}^{\mathrm{rat}}$ and $X_{C_{3}}\not\subset V_{3}^{\mathrm{rat}}$.
Remark 5.4.
Not every irreducible component of $V_{3}^{\mathrm{rat}}$ is of the form $\overline{X_{G}}$ for some $G$. The proof of Corollary 4.16 shows that $V_{3}^{\mathrm{rat}}\cap\overline{X_{C_{3}}}$ is a union of countably many (open, possibly singular) rational curves.

### v2

@@ 4.6. Identifying the complex torus $\mathrm{J}(\mathbb{V}^{\mu_{3}})$
Let $\mathcal{E}_{i}\rightarrow S$ be the relative elliptic curve with equation $y^{2}=x^{3}+{\rm disc}(f)^{i}$. The previous paragraph shows that $\mathrm{J}(\mathbb{W})$ is isomorphic to $\mathcal{E}_{i}({\rm\mathbb{C}})$ for some $i\in\{0,1,2,3,4,5\}$. We show that $i=1$, using our previous results on bielliptic Picard curves in [31]. Let $T\subset S$ be the closed subscheme where $b=0$, parametrizing even quartic polynomials $f=x^{4}+ax^{2}+c$. For $f\in T({\rm\mathbb{C}})$, the $\mu_{3}$-action on $\mathcal{C}_{f}$ extends to a $\mu_{6}$-action. A calculation shows ${\rm disc}|_{T}=16c(-a^{2}+4c)^{2}$. Applying the singular cohomology realization functor to [31, Theorem 5.1], the relative complex torus $\mathrm{J}((\mathbb{V}|_{T({\rm\mathbb{C}})})^{\mu_{6}})\rightarrow T({\rm\mathbb{C}})$ is isogenous to $\mathcal{E}_{1}({\rm\mathbb{C}})|_{T({\rm\mathbb{C}})}\rightarrow T({\rm\mathbb{C}})$. Since $\mathrm{J}((\mathbb{V}|_{T({\rm\mathbb{C}})})^{\mu_{6}})$ is a subtorus of $\mathrm{J}((\mathbb{V}|_{T({\rm\mathbb{C}})})^{\mu_{3}})$ of the same dimension, they must be equal, hence $\mathrm{J}(\mathbb{W})|_{T({\rm\mathbb{C}})}$ is isogenous to $\mathcal{E}_{1}({\rm\mathbb{C}})|_{T({\rm\mathbb{C}})}$ over $T({\rm\mathbb{C}})$. On the other hand, let $i\in\{0,1,2,3,4,5\}$ be such that $\mathrm{J}(\mathbb{W})\simeq\mathcal{E}_{i}({\rm\mathbb{C}})$. Then $\mathrm{J}(\mathbb{W})|_{T({\rm\mathbb{C}})}\simeq\mathcal{E}_{i}({\rm\mathbb{C}})|_{T({\rm\mathbb{C}})}$, hence $\mathcal{E}_{1}({\rm\mathbb{C}})|_{T({\rm\mathbb{C}})}$ is isogenous to $\mathcal{E}_{i}({\rm\mathbb{C}})|_{T({\rm\mathbb{C}})}$.
@@
Given $s=(a,b,c)\in S({\rm\mathbb{C}})$, consider the closed subscheme $T=\{(at,bt,ct^{2})\colon t\in{\rm\mathbb{A}}^{1}_{{\rm\mathbb{C}}}\}\cap S$ of $S$ and the restriction $\pi\colon\mathcal{E}|_{T}\rightarrow T$. The variety $\mathcal{E}|_{T}$ is an open subscheme of an elliptic surface with Weierstrass equation $y^{2}=x^{3}+t^{4}g(t)$, where $g(t)\in{\rm\mathbb{C}}[t]$ has degree $\leq 2$, using the formula (4.1). There exists a dense open $U\subset S$ such that for all $s\in U({\rm\mathbb{C}})$, $g(t)$ has two distinct nonzero roots. Let $s\in U({\rm\mathbb{C}})$. Using Tate’s algorithm, we see that the elliptic surface has a singular fiber above $t=0$ with Kodaira type $\text{\rm IV}^{*}$, singular fibers above the roots of $g(t)$ with Kodaira type II, and is smooth above $t=\infty$. The presence of a fiber of type II implies that $\mathcal{E}|_{T}\rightarrow T$ has no torsion sections [48, Lemma 7.8]. Moreover the Shioda–Tate formula [48, Theorem 6.3, Proposition 6.6 and §8.8] shows that the group of sections of $\mathcal{E}|_{T}\rightarrow T$ is free of rank $2$. Since $\mathcal{E}|_{T}$ receives an $\mathcal{O}$-action, its group of sections is free of rank $1$ over $\mathcal{O}$. We will now show that these facts can be used to prove the claims of the proposition by varying $s$ in $U({\rm\mathbb{C}})$.
@@ 4.7. Proof of the vanishing criterion in the Chow group
We claim that $\sigma$ is not a torsion section. If it were torsion, then ${\rm AJ}(\kappa_{\mathcal{C}})$ would be a torsion section of $\mathrm{J}(\mathbb{V}^{\mu_{3}})$, hence, by Proposition 4.12, $\kappa_{f}$ would be torsion for every $f\in S({\rm\mathbb{C}})$. This is not the case, since $\kappa_{f}$ is of infinite order if $f=x^{4}+x^{2}+1$ by [31, Corollary 2.9]. We conclude that $\sigma$ is not a torsion section.
@@
4.8. Analyzing the torsion locus of $P$
Theorem 4.2 shows that there are infinitely many plane quartic curves over ${\rm\mathbb{Q}}$ with torsion Ceresa cycle, since we may take $a=c=-12$. In fact, we can find explicit families of torsion Ceresa cycles over $\overline{{\rm\mathbb{Q}}}$ of arbitrarily large order by explicitly computing the torsion locus of the section $P$ of $\mathcal{E}\rightarrow S$.
Recall that $S\subset{\rm\mathbb{A}}^{3}_{{\rm\mathbb{C}}}$ is the parameter space of polynomials $f(x)=x^{4}+ax^{2}+bx+c$ of nonzero discriminant. For $(I,J)\in\mathbb{C}^{2}$ such that $4I^{3}-J^{2}\neq 0$, let $X_{I,J}\subset S$ be the closed subscheme of elements $f$ satisfying $(I(f),J(f))=(I,J)$, and let $E_{I,J}$ be the elliptic curve $y^{2}=x^{3}-Ix/3-J/27$. Let $E_{I,J}^{\circ}\subset E_{I,J}$ be the complement of the origin.
Lemma 4.16.
For $(I,J)\in{\rm\mathbb{C}}^{2}$ such that $4I^{3}-J^{2}\neq 0$, the map $f\mapsto(-2a/3,b)$ is an isomorphism $X_{I,J}\simeq E_{I,J}^{\circ}$.
@@ 4.8. Analyzing the torsion locus of $P$
Indeed, a calculation shows that an inverse is given by
(4.3) $(\alpha,\beta)\mapsto f(x)=x^{4}-3\alpha x^{2}/2+\beta x+(I/12-3\alpha^{2}/16).$
We now describe the locus of $f\in S$ where the specialization $P_{f}=(I(f),J(f))$ of the section $P$ of $\mathcal{E}\rightarrow S$ is torsion. First we consider the restriction of this locus to the closed subscheme $S_{1}\subset S$ where ${\rm disc}(f)=1$, so that $\mathcal{E}\rightarrow S$ restricts to the trivial family with fiber the elliptic curve $E_{0}\colon y^{2}=4x^{3}-27$. Let $E_{0}^{\circ}\subset E_{0}$ be the complement of the origin. If $(I,J)\in E_{0}^{\circ}({\rm\mathbb{C}})$, then $X_{I,J}$ is a closed subscheme of $S_{1}$. For $G$ a group, write $G_{tors}\subset G$ for the subset of torsion elements. Then we tautologically have
$\{f\in S_{1}(k)\colon P_{f}\in E_{0}({\rm\mathbb{C}})_{tors}\}=\bigsqcup_{(I,J)\in E_{0}({\rm\mathbb{C}})_{tors}}X_{I,J}({\rm\mathbb{C}}).$
Define a ${\rm\mathbb{G}}_{m}$-action on $S$ via the formula $\lambda\cdot(a,b,c)=(\lambda^{2}a,\lambda^{3}b,\lambda^{4}c)$. The polynomials $I(f),J(f)$ and ${\rm disc}(f)$ are then homogenous of degree $4,6,12$. Using the ${\rm\mathbb{G}}_{m}$-action on $S$, we can form the (coarse) quotient $\mathcal{M}_{Pic}=S/{\rm\mathbb{G}}_{m}$, an open subscheme of weighted projective space $\mathbb{P}(2,3,4)$.
Lemma 4.17.
The variety $\mathcal{M}_{Pic}$ is the coarse moduli space of Picard curves over ${\rm\mathbb{C}}$.
@@
See [26, p 15, Proposition 2.3]. ∎
The inclusion $S_{1}\subset S$ determines an isomorphism $S_{1}/\mu_{12}\simeq\mathcal{M}_{Pic}$, where $\mu_{12}\subset{\rm\mathbb{G}}_{m}$ acts via restriction on $S_{1}$. Write $\pi\colon S\rightarrow\mathcal{M}_{Pic}$ and $\pi_{1}\colon S_{1}\rightarrow\mathcal{M}_{Pic}$ for the quotient maps.
Let $V\subset\mathcal{M}_{Pic}$ be the image of the torsion locus in $S$ of the section $P$ of $\mathcal{E}\rightarrow S$. Since this torsion locus in ${\rm\mathbb{G}}_{m}$-invariant, $V$ also equals the image of the torsion locus of $P_{f}\in E_{0}$ under $\pi_{1}$. Therefore $V$ equals $\cup\pi_{1}(X_{I,J})$, where $(I,J)$ ranges over torsion points of $E_{0}^{\circ}$. If $\zeta\in\mu_{12}$, then $\zeta\cdot X_{I,J}=X_{\zeta^{4}I,\zeta^{6}J}$, so $\mu_{2}\subset\mu_{12}$ preserves $X_{I,J}$. Under the bijection $X_{I,J}\simeq E_{I,J}^{\circ}$, $-1$ acts as $(\alpha,\beta)\mapsto(\alpha,-\beta)$ on $E_{I,J}^{\circ}$. Therefore the quotient $X_{I,J}/\mu_{2}$ is isomorphic, via the $a$-coordinate, to ${\rm\mathbb{A}}^{1}$. It follows that for each torsion point $(I,J)$, the restriction $X_{I,J}\hookrightarrow S_{1}\rightarrow\mathcal{M}_{Pic}$ factors through a map $\varphi_{(I,J)}\colon X_{I,J}/\mu_{2}={\rm\mathbb{A}}^{1}\rightarrow\mathcal{M}_{Pic}$, and $V$ is the union of the images of $\varphi_{(I,J)}$. If $(I,J)=(\zeta^{4}I^{\prime},\zeta^{6}J^{\prime})$ for some $\zeta\in\mu_{12}$, then the images of $\varphi_{(I,J)}$ are equal. Otherwise, they are disjoint. We conclude:
Proposition 4.18.
We have
$V=\bigcup_{(I,J)\in E_{0}(\mathbb{C})_{tors}}\text{Image}(\varphi_{(I,J)}\colon{\rm\mathbb{A}}^{1}\rightarrow\mathcal{M}_{Pic}).$
The images of two maps $\varphi_{(I,J)}$ are either equal or disjoint. The maps $\varphi_{(I,J)}$ are quasi-finite.
We can identify $\varphi_{(I,J)}(t)$ explicitly, at least when $b\neq 0$. For $t\in{\rm\mathbb{C}}$, let $g_{(I,J)}(t)=t^{3}-It/3-J/27$, let $(\alpha_{t},\beta_{t})=(tg_{(I,J)}(t),g_{(I,J)}(t)^{2})$ and let $f_{(I,J),t}(x)=x^{4}-3\alpha_{t}x^{2}/2+\beta_{t}x+(I/12-3\alpha_{t}^{2}/16)$.
Proposition 4.19.
In the above notation, assume $g_{(I,J)}(t)\neq 0$. Then $\varphi_{(I,J)}(t)=\pi(f_{(I,J),t})$.
Proof.
Let $\tilde{f}$ be an element of $X_{I,J}$ which under the isomorphism $X_{I,J}\simeq E_{I,J}^{\circ}$ corresponds to an element $(\tilde{\alpha},\tilde{\beta})$ with $\tilde{\alpha}=t$. By definition, $\varphi_{(I,J)}(t)=\pi_{1}(\tilde{f})$. We have $\tilde{\beta}^{2}=g_{(I,J)}(t)\neq 0$. A calculation shows that $f_{(I,J),t}=\tilde{\beta}\cdot\tilde{f}$, so $f_{(I,J),t}$ and $\tilde{f}$ have the same image in $\mathcal{M}_{Pic}$, as claimed. ∎
Proposition 4.19 gives explicit one-parameter families of Picard curves with equation $C_{t}\colon y^{3}=f_{(I,J),t}(x)$ whose Ceresa cycle is torsion. The explicit description shows that when $(I,J)\in k^{2}$, where $k$ is a subfield of $\mathbb{C}$, then the family $C_{t}$ is defined over $k$. Since all torsion points of $E_{0}$ are defined over ${\rm\overline{\mathbb{Q}}}$, so are the families $C_{t}$.
We can now prove Theorem D.
Proof of Theorem D.
  1. $(i)$
    This follows from Theorems 4.2 and Proposition 4.18.
  2. $(ii)$
    For every $d\geq 1$, there exists $N=N(d)\geq 1$ such that for every Picard curve $C_{f}$ over a number field $k$ of degree $d$, the order of $\kappa_{f}$ in ${\rm CH}_{1}(J_{f,\bar{k}})$ is either infinite or less than $N$. Indeed, this follows from Theorem 4.2 and the uniform bound (depending only on $d$) on the order of a $k$-rational torsion point on an elliptic curve $y^{2}=x^{3}+D$ over any number field $k$ of degree $d$.
5. Ceresa vanishing loci in genus $3$
@@
These vanishing loci have the following basic properties: $V_{g}^{\mathrm{rat}}\subset V_{g}^{\mathrm{alg}}\subset\mathcal{M}_{g}$; the hyperelliptic locus is contained in $V_{g}^{\mathrm{rat}}$; and $V_{g}^{\mathrm{alg}}\neq\mathcal{M}_{g}$ by Ceresa’s famous result [14]. It would be interesting to obtain further information about the components of $V_{g}^{\mathrm{rat}}$ and $V_{g}^{\mathrm{alg}}$. We end our paper by determining the automorphism group strata in $\mathcal{M}_{3}$ that are contained in $V_{3}^{\mathrm{rat}}$ or $V_{3}^{\mathrm{alg}}$. So let $g=3$ and consider the open subscheme $\mathcal{M}_{3}^{\mathrm{nh}}\subset\mathcal{M}_{3}$ of non-hyperelliptic curves. There is a stratification $\mathcal{M}_{3}^{\mathrm{nh}}=\sqcup X_{G}$ into locally closed subvarieties such that a non-hyperelliptic curve $C$ over ${\rm\mathbb{C}}$ belongs to $X_{G}$ if and only if ${\rm Aut}(C)\simeq G$. It turns out that $X_{G}$ is irreducible and the closure of $X_{G}$ is a union of other strata. We refer to [35, §2.2] and references therein for a complete description of the loci $X_{G}$ and the closure relations between them. We reproduce here a diagram capturing these closure relations:
@@
For $n\in\{16,48,96\}$, the symbol $G_{n}$ the group of order $n$ and GAP label $(16,13),(48,33),$ and $(96,64)$ respectively [25]. See [35, Table 2] for models for a generic plane quartic in $X_{G}$. We make explicit the strata that are relevant for us: $\overline{X_{C_{3}}}$ is the locus of Picard curves studied in §4; $\overline{X_{C_{6}}}$ is the locus of bielliptic Picard curves studied in [32]; and the zero-dimensional strata each consist of a single automorphism-maximal curve with equation
@@
We end with the proof of Theorem E, which we restate for convenience:
@@
    Our analysis of Picard curves (Theorem C) shows that $X_{C_{3}}\subset V_{3}^{\mathrm{rat}}$, so $X_{G}\subset V_{3}^{\mathrm{rat}}$ for $G=C_{6},C_{9}$ and $G_{48}$ as well. On the other hand, the Ceresa cycle of the Fermat quartic and Klein quartic are known to be of infinite order in the Griffiths group; see [11, Theorem (4.1)] for the former and [28, §4] for the latter. By our observation, this means that $X_{G}\not\subset V_{3}^{\mathrm{rat}}$ for every stratum whose closure contains one of these curves. Since every stratum not contained in $\overline{X_{C_{3}}}$ has this property, we conclude the proof.
@@
    Since $V_{3}^{\mathrm{rat}}\subset V_{3}^{\mathrm{alg}}$, Part $(1)$ implies that $X_{G}\subset V_{3}^{\mathrm{rat}}$ only if $G=C_{3},C_{6},C_{9}$ or $G_{48}$. The criterion of Theorem A applies to the curves in $X_{C_{9}}$ and $X_{G_{48}}$ (see Example 3.1), so $V_{3}^{\mathrm{rat}}$ contains these strata. On the other hand, there exist curves in $X_{C_{6}}$ with nonvanishing $\kappa(C)$, by [31, Corollary 1.2]. So $X_{C_{6}}\not\subset V_{3}^{\mathrm{rat}}$ and $X_{C_{3}}\not\subset V_{3}^{\mathrm{rat}}$.